OpenStudy (anonymous):

The stem-and-leaf plot represents the average monthly high temperatures in Tampa, Florida. What is the median of the temperatures? Round to the nearest tenth. Steam Leaf ---- ------ 6 9 7 1 2 6 7 8 1 4 7 9 9 9 0 0 A. 81.0 B. 81.5 C. 82.0 D. 82.5

OpenStudy (anonymous):

There are 12 data value ranging from 69 up to 90 The median is the middle data value. But there are 2 data values at the middle: 81 and 84 The median is (81+84)/2 = 165/2 = 82.5